# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v2

@@ III.1 Qualitative analysis
${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{n}^{A}={\boldsymbol{\phi}}_{n}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
Here, we have chosen one of the L points at ${\bm{k}}_{n}$ as a representative one since the four L points are equivalent even when $\bm{h}$ is finite. For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
@@
Second, we note that the single-$\bm{q}$ free energy $F_{L}$ for the L point order parameter is isotropic. This means there is no anisotropy at the transition temperature $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ at ${\bm{H}}={\bf 0}$. This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$. When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus the third and fourth order terms in $F_{\Gamma}$ can be irrelevant. In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
@@
Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$ and low $T$, (c) small ${\bm{H}}$ and high $T$, and (d) large ${\bm{H}}$. (e) large ${\bm{H}}$. Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$). Filled red circles represent the directions $\theta_{L}$ given by Eq. (19) and open blue circles do $\theta_{\rm 0}$. A configuration for large ${\bm{H}}$ in the phase II in (c) is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
@@
Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$. The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crossovers toward one for $\bm{H}=0$. See Fig. 2(c). When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta F_{4}$ plays an important role. Thus, ${\boldsymbol{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig. 2(d). This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$. By the similar analysis to Eqs. (21)-(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig. 2(e)] is given as
@@ III.2 Phase diagram
Figures 3(a)–3(c) show the $T$-$H$ phase diagram for $\bm{H}\parallel[001]$, where the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$ , and (c) $b=0.01$ with $c=0.5$ fixed. As discussed in the previous section, the phase III is stabilized under high magnetic fields. This phase, however, is suppressed when $b$ becomes larger. We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$. The reason why the phase I cannot appear at the lower $T$ is understood by seeing Eq. (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
@@
For $\bm{H}\parallel[110]$, the situation becomes quite simpler. Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$. As discussed in the last part in the previous section, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively. For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on a (111) plane, while antiferroic between the neighboring (111) plane as shown in different colors in Fig. 4. When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=-1$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$. Color map represents the L point order parameter angle $\theta_{\rm L}$. (b) $\theta$$H$ phase diagram from [001]($\theta=0$) to [110] ($\theta=\pi/2$).
@@ IV.1 Free energy for symmetric states
In this section, we consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
@@
where $\bm{r}$ is labeled by the A sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same for simplicity (symmetric states). In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression. Thus, types of order realized below the second-order transition at $T=T_{c}$ are determined by the fourth-order terms.
@@
$F_{\rm L}^{2(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
@@
$F_{\rm L}^{4(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}$
$+$ $\frac{c^{\prime\prime\prime}}{64}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\Bigg{]}.$ (27)
@@
Note that the $c^{\prime\prime\prime}$ term can be negative, e.g., when $\theta_{1}=\theta_{2}=\theta_{3}=\theta_{4}+\pi$. This means that when $c^{\prime\prime\prime}$ is large, such quadruple-$\bm{q}$ orders are favored. The lowest free energy under the constraint (8) is $F_{\rm L}^{4(4)}=\tfrac{c}{4}\sum_{\alpha}\Phi_{\alpha}^{4}$ with $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(\theta,\theta,\theta^{\prime},\theta^{\prime}+\pi)$, where $\theta$ and $\theta^{\prime}$ are arbitrary and those with any two $\theta_{n}$’s permutations are equivalent ones. The unit cell for the quadrupole-$\bm{q}$ configurations contains eight A-sublattice sites and thus sixteen sites in total. See also Fig. 9 in Appendix A.
In Eqs. (26) and (27), $\theta_{n}$ dependence appears only in $c^{\prime\prime}$ and $c^{\prime\prime\prime}$ terms. Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane for (a) $c^{\prime}>2c$ and (b) $c^{\prime}<2c$ with $c>0$. For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from the expressions (26) and (27). See Fig. 6(a). The conditions given by Eq. (8) just correspond to the phase boundary for $c^{\prime\prime},c^{\prime\prime\prime}>0$ and $c^{\prime}=2c$, where all the single-$\bm{q}$, double-$\bm{q}$, and quadruple-$\bm{q}$ are degenerate. Note that away from the phase boundaries between quadruple-$\bm{q}$ and double- or single-$\bm{q}$ phases, the direction of the $\bm{\phi}_{n}$ in the quadruple-$\bm{q}$ states are pinned to $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$(0,0,0,0)$0$(0,0,0,0)( 0 , 0 , 0 , 0 ), $(0,0,\pi,\pi)$, $(0,\pi,\pi,0)$, $(0,\pi,0,\pi)$ for $c^{\prime\prime\prime}<0$, while $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$$(0,0,0,\pi)$, $(0,0,\pi,0)$, $(0,\pi,0,0)$, $(\pi,0,0,0)$ for $c^{\prime\prime\prime}>0$. Here, global sign changes $\bm{\phi}^{\alpha}\to-\bm{\phi}^{\alpha}$: $\theta_{n}^{\alpha}\to\theta_{n}^{\alpha}+\pi$, have not been included since they are also realized by trivial translations.
Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane. (a) $c^{\prime}>2c$, where the phase boundary between the single-$\bm{q}$(1q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2-(c^{\prime}-2c)$. (b) $c^{\prime}<2c$, where the phase boundary between the double-$\bm{q}$(2q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2+(2c-c^{\prime})/3$.
@@
So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters. Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$. In the lowest orders, the cubic terms in Eq. (5) are the most important. Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where ${\bm{k}}$ is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$. The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$. As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$. See Eq. (38). Furthermore, nontrivial couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$. They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
@@
The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable. However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$. This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ with $\theta_{n}-\theta_{m}=\pm\pi/2$. See the coupling in Eq. (13). As we will see in Sec. IV.2, the nontrivial terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.

### v3

@@ III.1 Qualitative analysis
${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{1}^{A}={\boldsymbol{\phi}}_{1}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
Here, we have chosen one of the L points at ${\bm{k}}_{1}$ as a representative one since the other three fields at ${\bm{k}}_{n}(n=2,3,4)$ with ${\boldsymbol{\phi}}_{n}^{A}=-{\boldsymbol{\phi}}_{n}^{B}$ are equivalent and this is also the case even when $\bm{H}$ is finite. See Table 1. For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
@@
Second, we note that the single-$\bm{q}$ free energy $F_{\rm L}$ is isotropic. This means that the transition temperatures for $u_{\rm L}$ and $v_{\rm L}$ are the same at ${\bm{H}}={\bf 0}$ and $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$. When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus, the third- and fourth-order terms in $F_{\Gamma}$ can be irrelevant. In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
@@
Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$, (c) small ${\bm{H}}$ and low $T$, and (d) large ${\bm{H}}$. Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$). Filled red circles represent the directions $\theta_{\rm L}$ given by Eq. (19) and open blue circles do $\theta_{\rm 0}$. A configuration for large ${\bm{H}}$ in the phase II is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
@@
Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$. The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crossovers toward one for $\bm{H}=0$. See Fig. 2(c). When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta F_{4}$ plays an important role. Thus, ${\boldsymbol{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig. 2(d). This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$. See Eq. (18). By the similar analysis to Eqs. (21)-(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig. 2(e)] is given as
@@ III.2 Phase diagram
Figures 3(a)–3(c) show the $T$-$H$ phase diagram for $\bm{H}\parallel[001]$, where the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$ , and (c) $b=0.01$ with $c=0.5$ fixed. As discussed in the previous section, the phase III is stabilized under high magnetic fields. This phase, however, is suppressed when $b$ becomes larger. We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$. The reason why the phase I cannot appear at the lower $T$ is understood by examining Eq. (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
@@
For $\bm{H}\parallel[110]$, the situation becomes quite simpler. Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$. As discussed in the last part in Sec. III.1, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively. For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on the (111) plane, while antiferroic between the neighboring (111) planes as shown in different colors in Fig. 4. When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$. Color map represents the L point order parameter angle $\theta_{\rm L}$. Single-$\bm{q}$ order with ${\bm{k}}={\bm{k}}_{1}$ for (b) the phase I and (c) the phase IV.
@@ IV.1 Free energy for symmetric states
We consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
@@
where $\bm{r}$ is labeled by the $A$ sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same $\Phi_{\alpha}/\sqrt{\mathcal{N}}(>0)$ for simplicity (symmetric states). For $n=1$ $\bm{\phi}^{A}_{1}=\bm{\phi}^{B}_{1}$, while, for $n=2,3$ and $4$, $\bm{\phi}^{A}_{n}=-\bm{\phi}^{B}_{n}$. As mentioned before, this is due to our unit cell convention. In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression. Thus, types of order realized below the second-order transition at $T=T_{c}$ are determined by the fourth-order terms.
@@
$F_{\rm L}^{2\bm{q}(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
@@
$F_{\rm L}^{4\bm{q}(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}\Bigg{]}.$ (27)
@@
Note that the $c^{\prime\prime\prime}$ terms cancel when summing up the terms for $\alpha=A$ and $B$, since $\phi^{A}_{n}=(-1)^{p_{n}}\phi^{B}_{n}$. The minimum value of $F_{\rm L}^{4\bm{q}(4)}$ for $c^{\prime\prime}>0$ is realized with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=2$, where two orthogonal pairs with the relative angle between the two is arbitrary, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta^{\alpha}+\pi/2,\theta^{\alpha},\varphi^{\alpha}+\pi/2,\varphi^{\alpha})$. For $c^{\prime\prime}<0$, the configurations for the minimum free energy is, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta,\theta,\theta,\theta)$ with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=6$. The unit cell for the quadrupole-$\bm{q}$ configurations contains eight $A$-sublattice sites with sixteen sites in total. See also Fig. 9 in Appendix A.
Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime}$$c^{\prime}$ plane for $c>0$. Note that there is no $c^{\prime\prime\prime}$ dependence. For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from Eqs. (26) and (27). The conditions given by Eq. (8) just correspond to the point $(c^{\prime}/c,c^{\prime\prime}/c)=(2,4)$, where the single-$\bm{q}$ and double-$\bm{q}$ are degenerate. As long as the local fourth-order terms are dominant, quadrupole-$\bm{q}$ orders are unlikely. Thus, the $c^{\prime}$ term is important when one compares the free energy for the single- and double-$\bm{q}$ orders.
Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime}$$c^{\prime\prime}$ plane for $c>0$. Filled circle indicate the bare parameter set $c^{\prime}=2c$ and $c^{\prime\prime}=4c$ given by Eq. (8). Region for $c^{\prime}+c^{\prime\prime}<-2c$ is unphysical since the fourth-order term becomes negative. For quadruple-$\bm{q}$ states, their representative configurations for $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})$ are indicated.
@@
So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters. Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$. In the lowest orders, the cubic terms in Eq. (5) are the most important. Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where ${\bm{k}}$ is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$. The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}\equiv\bf{0}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$. As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$. See Eq. (38). Furthermore, the fifth-order couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$. They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
@@
The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable. However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$. This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ and $\theta_{n}-\theta_{m}=\pm\pi/2$. See the coupling in Eq. (13). As we will see in Sec. IV.2, the fifth-order terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.